OpenStudy (anonymous):

help me solve this. lim 2x^3 +7x/3sin^2x x - 0

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

\[\lim_{x\to0}\frac{2x^3+7x}{3\sin^2(x)}\]??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then what will i do

OpenStudy (anonymous):

probably has no limit

OpenStudy (anonymous):

how it happen

OpenStudy (anonymous):

the lim of thatequation approaches 0

OpenStudy (anonymous):

what tools can you use? can you use l'hopital's rule?

OpenStudy (anonymous):

and you are sure it is \(x^3\) up top right?

OpenStudy (anonymous):

YES IT IS

OpenStudy (anonymous):

ok do you know l'hopital's rule?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

oh what are you supposed to use to solve this then?

OpenStudy (anonymous):

substitute 0 to the given equation?

OpenStudy (anonymous):

lol no that only tells you that you get \(\frac{0}{0}\) which is not a number just means you need to do more work

OpenStudy (anonymous):

how. what will i use

OpenStudy (anonymous):

i take it this is a calculus course right?

OpenStudy (anonymous):

yes it is.

OpenStudy (anonymous):

but you did not get to L'Hopital's rule yet we can try some algebra

OpenStudy (anonymous):

\[\frac{2x^3+7x}{3\sin^2(x)}=\frac{x}{3\sin(x)}\times \frac{2x^2+7}{\sin(x)}\]

OpenStudy (anonymous):

yes i didnt

OpenStudy (anonymous):

\[\lim_{x\to 0}\frac{x}{3\sin(x)}=\frac{1}{3}\] since \[\lim_{x\to 0}\frac{\sin(x)}{x}=1\]

OpenStudy (anonymous):

how the 2x^2+7 and 3sin be cancel?

OpenStudy (anonymous):

so we have \[\frac{1}{3}\lim_{x\to 0}\frac{2x^2+7}{\sin(x)}\] to consider

OpenStudy (anonymous):

now if you replace \(x\) by \(0\) in \[\frac{2x^2+7}{\sin(x)}\] you do not get \(\frac{0}{0}\) but rather you get \(\frac{7}{0}\) this means it is not "undetermined" i.e. you have more work to do this means the limit does not exist

OpenStudy (anonymous):

so theres no answer but it couldnt happen?

OpenStudy (anonymous):

the same way \[\lim_{x\to 0}\frac{1}{x}\] is undefined

OpenStudy (anonymous):

yes. but it cant happen there must be an answer

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

a limit does not necessarily exist, so no, there need not be an answer